# OC4-ReID: Occluded Cloth-Changing Person Re-Identification

### v5

@@ Abstract
The study of Cloth-Changing Person Re-identification (CC-ReID) focuses on retrieving specific pedestrians when their clothing has changed, typically under the assumption that the entire pedestrian images are visible. Pedestrian images in real-world scenarios, however, are often partially obscured by obstacles, presenting a significant challenge to existing CC-ReID systems. In this paper, we introduce a more challenging task termed Occluded Cloth-Changing Person Re-identification (OC4-ReID), which simultaneously addresses two challenges of clothing changes and occlusions. Concretely, we construct two new datasets, Occ-LTCC and Occ-PRCC, based on original CC-ReID datasets to include random occlusions of key pedestrians components (e.g., head, torso). Moreover, a novel benchmark is proposed for OC4-ReID incorporating a Train-Test Micro Granularity Screening (T2MGS) module to mitigate the influence of occlusions and proposing a Part-Robust Triplet (PRT) loss for partial features learning. Comprehensive experiments on the proposed datasets, as well as on two CC-ReID benchmark datasets demonstrate the superior performance of proposed method against other state-of-the-art methods. The codes and datasets are available at: https://github.com/1024AILab/OC4-ReID.
@@ I Introduction
Cloth-Changing Person Re-identification (CC-ReID) aims to retrieve target pedestrians across instances where their clothing has changed [1, 2, 3]. To mitigate clothing bias, existing methods utilize deep neural networks to learn features that are less dependent on clothing, often assuming that the pedestrian images are minimally occluded [4, 5, 1, 6, 7]. In real-world application scenarios, pedestrian images captured by practical cameras, however, are often occluded. This occlusion reduces available feature information and can cause the model to make incorrect matches.
Fig. 1 illustrates the evaluation of classical CC-ReID architectures [8, 9] on both the standard CC-ReID dataset and our newly constructed OC4-ReID dataset, where parts of the pedestrian’s body are randomly occluded. Our experimental results indicate that occlusions into the CC-ReID dataset significantly impairs the model’s performance, with CAL [8] and AIM [9] reducing in Rank-1 by 9.87% and 10.80%, and in mAP by 12.62% and 13.32%, respectively.
Given the significant impact of occlusions on CC-ReID, we introduce a novel and more challenging task, termed Occluded Cloth-Changing Person Re-identification (OC4-ReID). This task aims to identify pedestrians while overcoming the dual challenges of occlusions and clothing changes.
Figure 1: We construct OC4-ReID dataset based on the CC-ReID dataset by introducing random occlusions on various human body parts (e.g., head, arms) . Fair training and testing results indicate that traditional CC-ReID methods (e.g., CAL [8] and AIM [9]) perform worse on the OC4-ReID dataset.
Particularly, OC4-ReID is critical for suspects tracking, as suspects often change clothes and utilize obstacles to hide their bodies during escape, which misleads conventional ReID tracking systems. Compared to CC-ReID, OC4-ReID poses two primary challenges: (i) occlusion leads to even less cloth-unrelated pedestrian discrimination information; (ii) blurred and low-quality occlusion regions further interfere with effective features extraction.
@@ II-A Occluded Person Re-identification (OCC-ReID)
II-B Cloth-changing Person Re-identification (CC-ReID)
CC-ReID considers a long-term retrieval situation where the pedestrian’s clothing changes. To tackle this, the researchers employed different approaches including GAN [1], 3D information [34], human contour [3], and adversarial loss [8]. Xu et al. [1] enhanced the model’s ability to learn cloth-unrelated features by reconstructing intra-class examples to minimize feature differences while generating adversarial dressing examples across classes. Chen et al. [34] introduced 3D Shape Learning (3DSL), leveraging 3D body reconstruction as an auxiliary task to extract texture-insensitive 3D shape features for CC-ReID. Cui et al. [3] achieved controlled reconstruction by introducing the cloth-unrelated, cloth-related, and pedestrian contour features. Gu et al. [8] introduced a cloth-based adversarial loss which first trains a cloth classifier to capture cloth-related features, then shifts focus to learning cloth-unrelated features in the second stage.
@@ V-A Datasets and Evaluation Metrics
We evaluate the performance of our method on two CC-ReID benchmark datasets PRCC [11] and LTCC [12], as well as the proposed Occ-PRCC and Occ-LTCC datasets, under the cloth-changing configuration. PRCC [11] is a large CC-ReID dataset consists of 33,698 pedestrian images from 221 identities, and each person wears the same clothes in cameras 1 and 2, and different clothes in cameras 1 and 3. LTCC [12] contains 17,119 images from 152 identities and 478 outfits, which is captured with 12 different views. Following the previous studies [8, 9], this paper use Rank@1 and mean Average Precision (mAP) for evaluation.
@@ V-B Implementation Details
TABLE I: Comparison with state-of-the-art methods on the standard CC-ReID datasets and OC4-ReID datasets.
@@ V-C Compare with State-of-the-art Methods
We compare the proposed model with three traditional ReID architectures (i.e. HACNN [22], PCB [23], and IANet [24]) and three classical open-source CC-ReID methods (i.e. RCSANet [25], CAL [8], AIM [9]) in Table I.
@@
Result on PRCC and LTCC: In addition, we compare our method with current state-of-the-art methods on standard CC-ReID datasets: PRCC [11] and LTCC [12]. On the LTCC dataset, the Rank-1 and mAP of our method under the cloth-changing scenario are 43.84% and 22.43%, outperforming the second-best method by 3.24% and 3.33%, respectively. Similarly, our method also achieve the optimal performance on PRCC dataset. The experimental results demonstrate that our method can also achieve excellent performance in a separate cloth-changing scenario.
@@ VI Conclusion
In this paper, we propose the task of cloth-changing person re-identification under occlusion conditions, termed OC4-ReID, for the first time. To support this new task, we construct two specialized datasets, Occ-PRCC and Occ-LTCC. To address the challenges presented in this task, we design a new benchmark, which includes a Train-Test Micro Granularity Screening (T2MGS) module. This module effectively mitigates occlusions while facilitating the extraction of fine-grained cloth-unrelated features. Furthermore, we propose a Part-Robust Triplet (PRT) loss, which supervises the partial features learning properly. Experimental results on both CC-ReID and OC4-ReID datasets demonstrate the effectiveness of our proposed method.

### v6

@@ Abstract
The study of Cloth-Changing Person Re-identification ($C$-ReID) focuses on retrieving specific pedestrians when their clothing has changed, typically under the assumption that the entire pedestrian images are visible. Pedestrian images in real-world scenarios, however, are often partially obscured by obstacles, presenting a significant challenge to existing $C$-ReID systems. In this paper, we introduce a more challenging task termed Occluded Cloth-Changing Person Re-identification (OC4-ReID), which simultaneously addresses two challenges of clothing changes and occlusions. Concretely, we construct two new datasets, Occ-LTCC and Occ-PRCC, based on original $C$-ReID datasets to include random occlusions of key pedestrians components (e.g., head, torso). Moreover, a novel benchmark is proposed for OC4-ReID incorporating a Train-Test Micro Granularity Screening (T2MGS) module to mitigate the influence of occlusions and proposing a Part-Robust Triplet (PRT) loss for partial features learning. Comprehensive experiments on the proposed datasets, as well as on two $C$-ReID benchmark datasets demonstrate the superior performance of proposed method against other state-of-the-art methods. The codes and datasets are available at: https://github.com/1024AILab/OC4-ReID.
@@ I Introduction
Cloth-Changing Person Re-identification ($C$-ReID) aims to retrieve target pedestrians across instances where their clothing has changed [1, 2, 3]. To mitigate clothing bias, existing methods utilize deep neural networks to learn features that are less dependent on clothing, often assuming that the pedestrian images are minimally occluded [4, 5, 1, 6, 7]. In real-world application scenarios, pedestrian images captured by practical cameras, however, are often occluded. This occlusion reduces available feature information and can cause the model to make incorrect matches.
Fig. 1 illustrates the evaluation of classical $C$-ReID architectures [8, 9] on both the standard $C$-ReID dataset and our newly constructed OC4-ReID dataset, where parts of the pedestrian’s body are randomly occluded. Our experimental results indicate that occlusions into the $C$-ReID dataset significantly impairs the model’s performance, with CAL [8] and AIM [9] reducing in Rank-1 by 9.87% and 10.80%, and in mAP by 12.62% and 13.32%, respectively.
Given the significant impact of occlusions on $C$-ReID, we introduce a novel and more challenging task, termed Occluded Cloth-Changing Person Re-identification (OC4-ReID). This task aims to identify pedestrians while overcoming the dual challenges of occlusions and clothing changes.
Figure 1: We construct OC4-ReID dataset based on the $C$-ReID dataset by introducing random occlusions on various human body parts (e.g., head, arms) . Fair training and testing results indicate that traditional $C$-ReID methods (e.g., CAL [8] and AIM [9]) perform worse on the OC4-ReID dataset.
Particularly, OC4-ReID is critical for suspects tracking, as suspects often change clothes and utilize obstacles to hide their bodies during escape, which misleads conventional ReID tracking systems. Compared to $C$-ReID, OC4-ReID poses two primary challenges: (i) occlusion leads to even less cloth-unrelated pedestrian discrimination information; (ii) blurred and low-quality occlusion regions further interfere with effective features extraction.
@@ II-A Occluded Person Re-identification (OCC-ReID)
II-B Cloth-changing Person Re-identification ($C$-ReID)
$C$-ReID considers a long-term retrieval situation where the pedestrian’s clothing changes. To tackle this, the researchers employed different approaches including GAN [1], 3D information [34], human contour [3], and adversarial loss [8]. Xu et al. [1] enhanced the model’s ability to learn cloth-unrelated features by reconstructing intra-class examples to minimize feature differences while generating adversarial dressing examples across classes. Chen et al. [34] introduced 3D Shape Learning (3DSL), leveraging 3D body reconstruction as an auxiliary task to extract texture-insensitive 3D shape features for $C$-ReID. Cui et al. [3] achieved controlled reconstruction by introducing the cloth-unrelated, cloth-related, and pedestrian contour features. Gu et al. [8] introduced a cloth-based adversarial loss which first trains a cloth classifier to capture cloth-related features, then shifts focus to learning cloth-unrelated features in the second stage.
@@ V-A Datasets and Evaluation Metrics
We evaluate the performance of our method on two $C$-ReID benchmark datasets PRCC [11] and LTCC [12], as well as the proposed Occ-PRCC and Occ-LTCC datasets, under the cloth-changing configuration. PRCC [11] is a large $C$-ReID dataset consists of 33,698 pedestrian images from 221 identities, and each person wears the same clothes in cameras 1 and 2, and different clothes in cameras 1 and 3. LTCC [12] contains 17,119 images from 152 identities and 478 outfits, which is captured with 12 different views. Following the previous studies [8, 9], this paper use Rank@1 and mean Average Precision (mAP) for evaluation.
@@ V-B Implementation Details
TABLE I: Comparison with state-of-the-art methods on the standard $C$-ReID datasets and OC4-ReID datasets.
@@ V-C Compare with State-of-the-art Methods
We compare the proposed model with three traditional ReID architectures (i.e. HACNN [22], PCB [23], and IANet [24]) and three classical open-source $C$-ReID methods (i.e. RCSANet [25], CAL [8], AIM [9]) in Table I.
@@
Result on PRCC and LTCC: In addition, we compare our method with current state-of-the-art methods on standard $C$-ReID datasets: PRCC [11] and LTCC [12]. On the LTCC dataset, the Rank-1 and mAP of our method under the cloth-changing scenario are 43.84% and 22.43%, outperforming the second-best method by 3.24% and 3.33%, respectively. Similarly, our method also achieve the optimal performance on PRCC dataset. The experimental results demonstrate that our method can also achieve excellent performance in a separate cloth-changing scenario.
@@ VI Conclusion
In this paper, we propose the task of cloth-changing person re-identification under occlusion conditions, termed OC4-ReID, for the first time. To support this new task, we construct two specialized datasets, Occ-PRCC and Occ-LTCC. To address the challenges presented in this task, we design a new benchmark, which includes a Train-Test Micro Granularity Screening (T2MGS) module. This module effectively mitigates occlusions while facilitating the extraction of fine-grained cloth-unrelated features. Furthermore, we propose a Part-Robust Triplet (PRT) loss, which supervises the partial features learning properly. Experimental results on both $C$-ReID and OC4-ReID datasets demonstrate the effectiveness of our proposed method.